100 mL of 0.1 M HCl are mixed with 100 mL of 0.25
Answer #1

V = 100 ml = 0.1 L

M = 0.1 M HCl

V2 = 100 ml = 0.1

M = 0.25 M of NOH

dT = 5

Hrxn = Q/n

Q = m*Cp*dT = (100+100)*4.18*5 = 4180 J

n = 0.1*1 = 0.1 mol

Hrxn = -Q/n = -4180/0.1 = -41800 J/mol = -41.8 kJ/mol